OpenStudy (anonymous):

x + 3y = 5 -x + 6y = 4 Solve the system of equations. A. x = 1, y = 2 B. x = 2, y = 1 C. x = 1, y = 1 D. x = 0, y = 2 @princeharryyy

OpenStudy (anonymous):

@uri @KlOwNlOvE

OpenStudy (anonymous):

@lala_

OpenStudy (anonymous):

lets use elimination on this being we can solve y almost right away. combine like terms in both equations. x will cancel out 9y=9 solve for y by dividing 9 by each side then multiply the first equation by -2 -2(x+3y=5) new systems will be -2x-6y=-10 -x+6y=4 Combine like terms y's will cancel out -3x=-6 divide both sides by -3 (2 negatives will make a positive)